# NVRC: Neural Video Representation Compression

### v1

@@ Abstract
Recent advances in implicit neural representation (INR)-based video coding have demonstrated its potential to compete with both conventional and other learning-based approaches. With INR methods, a neural network is trained to overfit a video sequence, with its parameters compressed to obtain a compact representation of the video content. However, although promising results have been achieved, the best INR-based methods are still out-performed by the latest standard codecs, such as VVC VTM, partially due to the simple model compression techniques employed. In this paper, rather than focusing on representation architectures as in many existing works, we propose a novel INR-based video compression framework, Neural Video Representation Compression (NVRC), targeting compression of the representation. Based on the novel entropy coding and quantization models proposed, NVRC, for the first time, is able to optimize an INR-based video codec in a fully end-to-end manner. To further minimize the additional bitrate overhead introduced by the entropy models, we have also proposed a new model compression framework for coding all the network, quantization and entropy model parameters hierarchically. Our experiments show that NVRC outperforms many conventional and learning-based benchmark codecs, with a 24% average coding gain over VVC VTM (Random Access) on the UVG dataset, measured in PSNR. As far as we are aware, this is the first time an INR-based video codec achieving such performance. The implementation of NVRC will be released at www.github.com.
@@ 1 Introduction
In recent years, learning-based video compression [35, 29, 31, 11, 25] has demonstrated its significant potential to compete with conventional video coding standards, with some recent contributions (e.g., DCVC-DC [31]) reported to outperform the latest MPEG standard codec, VVC VTM [8]. However, learning-based codecs are typically associated with high computational complexity, in particular at the decoder, which therefore limits their practical deployment. To address this, a new type of learned video codec has been proposed, based on implicit neural representation (INR) models [11, 25], where each INR instance is overfitted and compressed to represent a video sequence (or a video dataset). INR-based codecs enable much faster decoding speed compared to most the non-INR learning based coding methods, and do not require offline optimization due to its overfitting nature. Although they have shown promise, INR-based codecs are yet to compete with state-of-the-art conventional and learned video coding methods in terms of rate-distortion performance.
To enhance coding performance, it is noted that most recent INR-based video coding methods [11, 25] focus on improving network architectures but still perform simply model pruning, quantization and entropy coding to obtain compact representations. Moreover, these methods are not fully end-to-end optimized; for example, NeRV [11] and HiNeRV [25] are not trained with a rate-distortion objective but only perform fine-tuning with pruning and quantization applied. Although COOL-CHIC [28] and C3 [24] are almost end-to-end optimized, the rate consumed by their entropy model and decoder/synthesis networks does not contribute to the training process. In contrast, state-of-the-art non-INR learning-based codecs [29, 30] are typically end-to-end trained with advanced entropy models, and this contributes to their improved coding performance compared to INR-based methods.
@@
Figure 1: Comparison between the output from HiNeRV [25] and the proposed NVRC. The image is from UVG dataset (Jockey/ReadySetGo sequence) [40]
@@
We conducted an experiment to compare the proposed approach with state-of-the-art conventional and learning-based video codecs on the UVG [40] dataset. To enable a fair comparison, we use both the RGB444 (like other learned video codecs) and YUV420 (like standard video coding methods) configurations. The results demonstrate the effectiveness of NVRC, which achieved up to 24% and 50% BD-rate savings when compared to the latest MPEG standard codec, H.266/VVC VTM-20.0 (Random Access), and the state-of-the-art INR-based codec, HiNeRV [25], respectively. To our best knowledge, NVRC is the first INR-based video codec outperforming VVC VTM with such significant coding gains.
@@ 2.1 Learning-based video compression
Video compression is an important research topic that underpins the development of many video-related applications, such as video streaming, video conferencing, surveillance, and gaming. In the past three decades, multiple generations of video coding standards [54, 51, 8] have evolved by integrating advanced coding techniques. Recently, learning-based video compression emerged as an popular alternative due to its strong expressive power and the ability to be optimized in a data- and metric-driven manner. Neural networks can be combined with conventional codecs for performance enhancement [37, 57] or used to build end-to-end optimized frameworks. DVC [35] first proposed to replace all modules in conventional hybrid codecs using neural networks. Follow-up innovations include those improving motion estimation [33, 2, 21, 31], applying feature space conditional coding [22, 29], and adopting novel architectures such as normalizing flows [20], transformers [55, 39], etc. In addition to these architecture modifications, improvements to quantization-involved optimization have also been achieved to handle the non-differentiability caused by hard thresholding operations [24, 3, 16]. Moreover, several studies [15, 46, 23, 36, 34] have validated the effectiveness of adapting a model to an individual image of a video sequence in order to eliminate the amortization gap [56, 53]. Despite demonstrating impressive rate-distortion performance, with some recent advancements reporting outperformance of VVC, neural video codecs are generally too computationally intensive [43], thus limiting their adoption in practical applications.
@@ 2.2 INR-based video compression
Implicit neural representation (INR) [47] is an emerging paradigm for representing multimedia data, such as audios [50], images [12, 49], videos [11], and 3D scenes [41, 44]. This type of method exploits the mapping from the coordinate inputs to a high-dimensional feature space and aims to output the corresponding target data value at that location. Neural representation for videos (NeRV) [11] has been proposed to model the mapping from frame indices to video frames, showing competitive reconstruction performance with a very high decoding speed. When applied to video compression, the network parameters of these models are compressed through pruning, quantization, and entropy-penalization [17, 14, 58] to achieve high coding efficiency. The following contributions further investigated patch-wise [5], volume-wise [38], or spatial-temporal disentangled representations [32] to improve representational flexibility. There are also methods that explicitly model the volume-wise residual [38], frame-wise residual [60], or flow-based motion compensation [59, 27, 18, 28], to enable scalable encoding and representation of longer and more diverse videos. In addition to these index-based approaches, other work has exploited content-specific embeddings/feature grids to provide visual prior for the network. The embeddings may be associated with single [10] or multiple resolutions [27, 25, 24, 28]. Although they hold promise in terms of low decoding complexity and competitive performance, all these aforementioned INR-based video codecs are still outperformed by state-of-the-art conventional (e.g., VVC VTM [8]) and autoencoder-based [30, 31] video codecs.
@@ 3 Method
Figure 2 shows the proposed NVRC framework, which follows a workflow similar to existing INR-based video compression methods, such as [11, 25, 24], but with a more advanced model compression pipeline. It trains a neural representation for a given video(s) and utilizes model compression techniques to obtain the compact representation (with compressed network parameters) of the video(s). Specifically, in NVRC, for a target video sequence $V^{gt}$ with $T$ frames, height $H$, width $W$, and $C$ channels, i.e., $V^{gt}\in\mathbb{R}^{T\times H\times W\times C}$, a neural representation $F$ parameterized by $\theta$ is trained to map coordinates to pixel intensities such as RGB colors, in a patch-wise manner. This can be represented by:
@@
where $i,j,k$ is the patch coordinates. $V_{patch}\in\mathbb{R}^{T_{patch}\times H_{patch}\times W_{patch}\times C_{patch}}$ is the corresponding video patch, in which $0\leq i<\frac{W}{W_{patch}}$, $0\leq j<\frac{H}{H_{patch}}$ and $0\leq k<\frac{T}{T_{patch}}$. This formulation (the same as in [25]) generalizes different frameworks: when $(T_{patch},H_{patch},W_{patch})=(1,1,1)$, the neural network maps coordinates to individual pixels [47]; when $(T_{patch},H_{patch},W_{patch})=(1,H,W)$, the network maps coordinates to video frames [11].
As mentioned in Section 1, existing INR-based video codecs either split the training of the INR model and model compression [11, 25], or train the model with compression techniques applied, but not entirely in an end-to-end manner [24, 28]. Unlike these works, NVRC employs more advanced entropy models and allows real end-to-end optimization. Specifically, a neural representation $F$ contains a set of learnable parameters $\theta$ including feature grids parameters ($\theta_{grid}$) and network layer parameters ($\theta_{layer}$). To quantize and encode these parameters, quantization and entropy models are employed with learnable compression parameters $\phi=\{\phi_{quant},\phi_{em}\}$. $\phi$ is determined by the distribution of the representation parameters $\theta$ in a fine-grained manner, e.g., per-group quantization scales, and can be considered as side information in compression [6]. While these parameters do improve overall coding efficiency, the introduced overhead is not negligible. Therefore $\phi$ is also quantized in this work, denoted as $\hat{\phi}=\{\hat{\phi}_{quant},\hat{\phi}_{em}\}$, and entropy coded. Here the quantization and entropy coding are performed based on another set of learnable parameters $\psi=\{\psi_{quant},\psi_{em}\}$, and can be simply quantized into full/half precision as $\hat{\psi}=\{\hat{\psi}_{quant},\hat{\psi}_{em}\}$. This forms a hierarchical coding strategy for encoding these model and compression parameters $\theta$, $\phi$ and $\psi$, as illustrated in Figure 2.
@@ 3.1 Feature grid coding
Although employing feature grids [10, 25, 24, 28, 26] for neural representations improves both convergence rate and reconstruction quality, these typically rely on a large number of parameters, which could potentially challenge model compression techniques. To address this issue, related works utilize multi-resolution grids to improve parameter efficiency and/or perform entropy coding for the feature grid compression [10, 27, 25, 24, 28, 26].
@@
3D block partitioning. For a feature grid $z$ in $\theta_{grid}$, with $z\in\mathbb{R}^{T_{grid}\times H_{grid}\times W_{grid}\times C_{grid}}$, it is divided into blocks with a size of $T_{blk}\times H_{blk}\times W_{blk}\times C_{grid}$ (padding is applied if the grid size is not divisible) and produced $\frac{T_{grid}}{T_{blk}}\times\frac{H_{grid}}{H_{blk}}\times\frac{W_{grid}}{W_{blk}}$ blocks. For multi-scale grids [27, 25, 28, 24], the same block size is applied to partition the grids at each scale.
Quantization. With the partitioned blocks, a transformation is then applied before quantization. Here, the corresponding per-block quantization scales $\delta^{grid}_{blk}$ from $\hat{\phi}^{quant}$ are utilized, where $\delta^{grid}_{blk}\in\mathbb{R}^{\frac{T_{grid}}{T_{blk}}\times\frac{H_{grid}}{H_{blk}}\times\frac{W_{grid}}{W_{blk}}\times C_{grid}}$ is learnable, and all the scales in $\delta^{grid}_{blk}$ are shared by features in the same channel and in the same block in $z$. To achieve this, $\delta^{grid}_{blk}$ is first expanded to $\delta^{grid}$, which has the same shape as $z$, and such that:
$\delta^{grid}[t,h,w,c]=\delta^{grid}_{blk}[\lfloor\frac{t}{T_{blk}}\rfloor,\lfloor\frac{h}{H_{blk}}\rfloor,\lfloor\frac{w}{W_{blk}}\rfloor,c],$ (2)
where $0\leq t<T_{grid}$, $0\leq h<H_{grid}$, $0\leq w<W_{grid}$ and $0\leq c<C_{grid}$. In practice, the logarithm of $\delta^{grid}_{blk}$ is learned, instead of $\delta^{grid}_{blk}$, which ensures that the scales are non-negative.
Following [14, 38, 24, 58, 26], the scaling and quantization can then be computed by:
$\hat{z}^{s}=\lfloor z^{s}\rceil=\lfloor\frac{z}{\delta^{grid}}\rceil,$ (3)
@@
$\hat{z}$ $=\hat{z}^{s}\times\delta^{grid},$ (4)
@@
Context-based entropy model. Although entropy coding has been used for reducing the bit rate consumed by feature grids, many implementations are only based on simple entropy models and treat the grids as ordinary network parameters [10, 27, 25, 26]. A better solution is to exploit the spatial-temporal redundancy within the feature grids, due to the inter dependent nature of the features. COOL-CHIC [28] and C3 [24] utilize context-based model and achieve efficient feature grid encoding; however, these methods are not associated with optimal rate distortion performance due to their low complexity constraints, and the use of grid entropy models for INR-based video compression has not been fully explored in the literature.
To enhance the efficiency of feature grid coding, NVRC employs context-based Gaussian models with auto-regressive style encoding and decoding processes [42] to exploit spatial-temporal redundancy within feature grids. While auto-regressive style coding is sequential, in NVRC, the feature grids at different resolutions are coded independently. Moreover, as mentioned above, each grid is partitioned into many small 3D blocks, which are also coded in parallel. Thus, the context model in NVRC has a high degree of parallelism, which enables fast coding despite of reduced amount of available context. The context-based model employs 3D masked convolution [52, 42], with parameters $\omega^{context}$ from $\hat{\phi}_{em}$, which are applied to all blocks from the same feature grid, but are not shared between grids at different scales. The context model estimates the means $\mu^{grid}$ and scales $\sigma^{grid}$ for the Gaussian distribution in a per-feature manner and uses a coder such as the arithmetic coder to code the parameters, i.e. $\hat{z}^{s}$. Here, the estimated means $\mu^{grid}$ and scales $\sigma^{grid}$ will also be scaled by the corresponding quantization scale $\delta^{grid}$ before applying for encoding and decoding.

### v2

@@ Abstract
Recent advances in implicit neural representation (INR)-based video coding have demonstrated its potential to compete with both conventional and other learning-based approaches. With INR methods, a neural network is trained to overfit a video sequence, with its parameters compressed to obtain a compact representation of the video content. However, although promising results have been achieved, the best INR-based methods are still out-performed by the latest standard codecs, such as VVC VTM, partially due to the simple model compression techniques employed. In this paper, rather than focusing on representation architectures, which is a common focus in many existing works, we propose a novel INR-based video compression framework, Neural Video Representation Compression (NVRC) [^1] Project page: https://hmkx.github.io/nvrc/, targeting compression of the representation. Based on its novel quantization and entropy coding approaches, NVRC is the first framework capable of optimizing an INR-based video representation in a fully end-to-end manner for the rate-distortion trade-off. To further minimize the additional bitrate overhead introduced by the entropy models, NVRC also compresses all the network, quantization and entropy model parameters hierarchically. Our experiments show that NVRC outperforms many conventional and learning-based benchmark codecs, with a 23% average coding gain over VVC VTM (Random Access) on the UVG dataset, measured in PSNR. As far as we are aware, this is the first time an INR-based video codec achieving such performance.
@@ 1 Introduction
In recent years, learning-based video compression lu2019dvc; li2021deep; li2023neural; chen2021nerv; kwan2023hinerv has demonstrated its significant potential to compete with conventional video coding standards, with some recent contributions (e.g., DCVC-DC li2023neural) reported to outperform the latest MPEG standard codec, VVC VTM bross2021overview. However, learning-based codecs are typically associated with high computational complexity, in particular at the decoder, which therefore limits their practical deployment. To address this, a new type of learned video codec has been proposed, based on implicit neural representation (INR) models chen2021nerv; kwan2023hinerv, where each INR instance is overfitted and compressed to represent a video sequence (or a video dataset). INR-based codecs enable much faster decoding speed compared to most non-INR learning based coding methods, and do not require offline optimization due to its overfitting nature. Although they have shown promise, INR-based codecs are yet to compete with state-of-the-art conventional and learned video coding methods in terms of rate-distortion performance.
To enhance coding performance, it is noted that most recent INR-based video coding methods chen2021nerv; kwan2023hinerv focus on improving network architectures but still perform simply model pruning, quantization and entropy coding to obtain compact representations. Moreover, these methods are not fully end-to-end optimized; for example, NeRV chen2021nerv and HiNeRV kwan2023hinerv are not trained with a rate-distortion objective but only perform fine-tuning with pruning and quantization applied. Although COOL-CHIC leguay2024cool and C3 kim2023c3 are almost end-to-end optimized, the rate consumed by their entropy model and decoder/synthesis networks does not contribute to the training process. In contrast, state-of-the-art non-INR learning-based codecs li2021deep; li2022hybrid are typically end-to-end trained with advanced entropy models, and this contributes to their improved coding performance compared to INR-based methods.
@@
Figure 1: Comparison between the output from HiNeRV kwan2023hinerv and the proposed NVRC. The image is from UVG dataset (Jockey/ReadySetGo sequence) mercat2020uvg.
@@
We conducted experiments to compare the proposed approach with state-of-the-art conventional and learning-based video codecs on the UVG mercat2020uvg, MCL-JCV wang2016mcl and JVET-CTC Class B boyce2018jvet datasets. To enable a fair comparison, we use both the RGB444 (like most learned video codecs) and YUV420 (like standard video coding methods) configurations. The results demonstrate the effectiveness of NVRC, which achieved up to 23% and 50% BD-rate savings when compared to the latest MPEG standard codec, H.266/VVC VTM-20.0 (Random Access) vtm, and the state-of-the-art INR-based codec, HiNeRV kwan2023hinerv, respectively. To our best knowledge, NVRC is the first INR-based video codec outperforming VVC VTM with such significant coding gains.
@@ 2.1 Learning-based video compression
Video compression is an important research topic that underpins the development of many video-related applications, such as video streaming, video conferencing, surveillance, and gaming. In the past three decades, multiple generations of video coding standards wiegand2003overview; sullivan2012overview; bross2021overview have evolved by integrating advanced coding techniques. Recently, learning-based video compression emerged as an popular alternative due to its strong expressive power and the ability to be optimized in a data- and metric-driven manner. Neural networks can be combined with conventional codecs for performance enhancement ma2020mfrnet; zhang2020enhancing or used to build end-to-end optimized frameworks. DVC lu2019dvc first proposed to replace all modules in conventional codecs using neural networks. Follow-up innovations include those improving motion estimation lin2020m; agustsson2020scale; hu2022coarse; li2023neural, applying feature space conditional coding hu2021fvc; li2021deep, optimizing context modeling in terms of performance and efficiency he2021checkerboard; he2022elic; zhou2022riddle, and adopting novel architectures such as normalizing flows ho2022canf, transformers xiang2022mimt; mentzer2022vct, etc. In addition to these architecture modifications, improvements to quantization-involved optimization have also been achieved to handle the non-differentiability caused by hard thresholding operations kim2023c3; agustsson2020universally; guo2021soft. Moreover, several studies guo2020variable; shen2023dec; khani2021efficient; lv2023dynamic; lu2020content have validated the effectiveness of adapting a model to an individual image of a video sequence via iterative refinement to reduce the amortization gap yang2020improving; van2021overfitting and optimize bit allocation over a sequence of frames xu2023bit. Despite demonstrating impressive rate-distortion performance, with some recent advancements reporting outperformance of VVC bross2021overview, neural video codecs are generally too computationally intensive minnen2023advancing, thus limiting their adoption in practical applications.
@@ 2.2 INR-based video compression
Implicit neural representation (INR) sitzmann2020implicit is an emerging paradigm for representing multimedia data, such as audios su2022inras, images dupont2021coin; strumpler2022implicit, videos chen2021nerv, and 3D scenes mildenhall2021nerf; muller2022instant. This type of method exploits the mapping from the coordinate inputs to a high-dimensional feature space and aims to output the corresponding target data value at that location. Neural representation for videos (NeRV) chen2021nerv has been proposed to model the mapping from frame indices to video frames, showing competitive reconstruction performance with a very high decoding speed. When applied to video compression, the network parameters of these models are compressed through pruning, quantization, and entropy-penalization han2015deep; gomes2023video; guo2023compression; zhang2024boosting to achieve high coding efficiency. The following contributions further investigated patch-wise bai2023ps, volume-wise maiya2023nirvana, or spatial-temporal disentangled representations li2022nerv to improve representational flexibility. There are also methods that explicitly model the volume-wise residual maiya2023nirvana, frame-wise residual zhao2023dnerv, or flow-based motion compensation zhang2021implicit; lee2023ffnerv; he2023towards; leguay2024cool, to enable scalable encoding and representation of longer and more diverse videos. In addition to these index-based approaches, other work has exploited content-specific embeddings/feature grids to provide visual prior for the network. The embeddings may be associated with single chen2023hnerv or multiple resolutions lee2023ffnerv; kwan2023hinerv; kim2023c3; leguay2024cool. Although they hold promise in terms of low decoding complexity and competitive performance, all these aforementioned INR-based video codecs are still outperformed by state-of-the-art conventional (e.g., VVC VTM bross2021overview) and autoencoder-based li2022hybrid; li2023neural video codecs.
@@ 3 Method
Figure 2 shows the proposed NVRC framework, which follows a workflow similar to existing INR-based video compression methods, such as chen2021nerv; kwan2023hinerv; kim2023c3, but with a more advanced model compression pipeline. It trains a neural representation for a given video(s) and utilizes model compression techniques to obtain the compact representation (with compressed network parameters) of the video(s). Specifically, in NVRC, for a target video sequence $V^{gt}$ with $T$ frames, height $H$, width $W$, and $C$ channels, i.e., $V^{gt}\in\mathbb{R}^{T\times H\times W\times C}$, a neural representation $F$ parameterized by $\theta$ is trained to map coordinates to pixel intensities such as RGB colors, in a patch-wise manner. This can be represented by:
@@
where $i,j,k$ is the patch coordinates. $V_{patch}\in\mathbb{R}^{T_{patch}\times H_{patch}\times W_{patch}\times C_{patch}}$ is the corresponding video patch, in which $0\leq i<\frac{W}{W_{patch}}$, $0\leq j<\frac{H}{H_{patch}}$ and $0\leq k<\frac{T}{T_{patch}}$. This formulation (the same as in kwan2023hinerv) generalizes different frameworks: when $(T_{patch},H_{patch},W_{patch})=(1,1,1)$, the neural network maps coordinates to individual pixels sitzmann2020implicit; when $(T_{patch},H_{patch},W_{patch})=(1,H,W)$, the network maps coordinates to video frames chen2021nerv.
As mentioned in Section 1, existing INR-based video codecs either split the training of the INR model and model compression chen2021nerv; kwan2023hinerv, or train the model with compression techniques applied, but not entirely in an end-to-end manner kim2023c3; leguay2024cool. Unlike these works, NVRC employs more advanced entropy models and allows fully end-to-end optimization. Specifically, a neural representation $F$ contains a set of learnable parameters $\theta$ including feature grids parameters ($\theta_{grid}$) and network layer parameters ($\theta_{layer}$). To quantize and encode these parameters, quantization and entropy models are employed with learnable compression parameters $\phi=\{\phi_{quant},\phi_{em}\}$. $\phi$ is determined by the distribution of the representation parameters $\theta$ in a fine-grained manner, e.g., per-group quantization scales, and can be considered as side information in compression balle2018variational. While these parameters do improve overall coding efficiency, the introduced overhead is not negligible, particularly when the compression ratio of the representation parameters is high. Therefore $\phi$ is also quantized in this work, denoted as $\hat{\phi}=\{\hat{\phi}_{quant},\hat{\phi}_{em}\}$, and entropy coded. Here the quantization and entropy coding are performed based on another set of learnable parameters $\psi=\{\psi_{quant},\psi_{em}\}$, which can be simply quantized into full/half precision as $\hat{\psi}=\{\hat{\psi}_{quant},\hat{\psi}_{em}\}$. This forms a hierarchical coding strategy for encoding these model and compression parameters $\theta$, $\phi$ and $\psi$, as illustrated in Figure 2.
@@ 3.1 Feature grid coding
Although employing feature grids chen2023hnerv; kwan2023hinerv; kim2023c3; leguay2024cool; kwan2024immersive for neural representations improves both convergence rate and reconstruction quality, these typically rely on a large number of parameters, which could potentially challenge model compression techniques. To address this issue, related works utilize multi-resolution grids to improve parameter efficiency and/or perform entropy coding for the feature grid compression chen2023hnerv; lee2023ffnerv; kwan2023hinerv; kim2023c3; leguay2024cool; kwan2024immersive.
@@
3D block partitioning. For a feature grid $z$ in $\theta_{grid}$, with $z\in\mathbb{R}^{T_{grid}\times H_{grid}\times W_{grid}\times C_{grid}}$, it is divided into blocks with a size of $T_{blk}\times H_{blk}\times W_{blk}\times C_{grid}$ (padding is applied if the grid size is not divisible) and produced $\frac{T_{grid}}{T_{blk}}\times\frac{H_{grid}}{H_{blk}}\times\frac{W_{grid}}{W_{blk}}$ blocks. For multi-scale grids lee2023ffnerv; kwan2023hinerv; leguay2024cool; kim2023c3, the same block size is applied to partition the grids at each scale.
Quantization. With the partitioned blocks, a transformation is then applied before quantization. Here, the corresponding per-block quantization scales $\delta_{grid,blk}$ from $\hat{\phi}_{quant}$ are utilized, where $\delta_{grid,blk}\in\mathbb{R}^{\frac{T_{grid}}{T_{blk}}\times\frac{H_{grid}}{H_{blk}}\times\frac{W_{grid}}{W_{blk}}\times C_{grid}}$ is learnable, and all the scales in $\delta_{grid,blk}$ are shared by features in the same channel and in the same block in $z$. To achieve this, $\delta_{grid,blk}$ is first expanded to $\delta^{grid}$, which has the same shape as $z$, and such that:
$\delta_{grid}[t,h,w,c]=\delta_{grid,blk}[\lfloor\frac{t}{T_{blk}}\rfloor,\lfloor\frac{h}{H_{blk}}\rfloor,\lfloor\frac{w}{W_{blk}}\rfloor,c],$ (2)
where $0\leq t<T_{grid}$, $0\leq h<H_{grid}$, $0\leq w<W_{grid}$ and $0\leq c<C_{grid}$. In practice, the logarithm of $\delta_{grid,blk}$ is learned, instead of $\delta_{grid,blk}$, which ensures that the scales are non-negative.
Following gomes2023video; maiya2023nirvana; kim2023c3; zhang2024boosting; kwan2024immersive, the scaling and quantization can then be computed by:
$\hat{z}^{s}=\lfloor z^{s}\rceil=\lfloor\frac{z}{\delta_{grid}}\rceil,$ (3)
@@
$\hat{z}$ $=\hat{z}^{s}\times\delta_{grid},$ (4)
@@
Context-based entropy model. Although entropy coding has been used for reducing the bit rate consumed by feature grids, many implementations are only based on simple entropy models and treat the grids as ordinary network parameters chen2023hnerv; lee2023ffnerv; kwan2023hinerv; kwan2024immersive. A better solution is to exploit the spatial-temporal redundancy within the feature grids, due to the inter dependent nature of the features. COOL-CHIC leguay2024cool and C3 kim2023c3 utilize context-based model and achieve efficient feature grid encoding; however, these methods are not associated with optimal rate distortion performance due to their low complexity constraints, and the use of grid entropy models for INR-based video compression has not been fully explored in the literature.
To enhance the efficiency of feature grid coding, NVRC employs context-based Gaussian models with auto-regressive style encoding and decoding processes minnen18joint to exploit spatial-temporal redundancy within feature grids. While auto-regressive style coding is sequential, in NVRC, the feature grids at different resolutions are coded independently. Moreover, as mentioned above, each grid is partitioned into many small 3D blocks, which are also coded in parallel. Thus, the context model in NVRC has a high degree of parallelism, which enables fast coding despite of reduced amount of available context. The context-based model employs 3D masked convolution oord16conditional; minnen18joint, with parameters $\omega_{context}$ from $\hat{\phi}_{em}$, which are applied to all blocks from the same feature grid, but are not shared between grids at different scales. The context model estimates the means $\mu_{grid}$ and scales $\sigma_{grid}$ for the Gaussian distribution in a per-feature manner and uses a coder such as the arithmetic coder to code the parameters, i.e., $\hat{z}^{s}$. Here, the estimated means $\mu_{grid}$ and scales $\sigma_{grid}$ will also be scaled by the corresponding quantization scale $\delta_{grid}$ before applying for encoding and decoding.